# Jonathan Light1  Min Cai2  Weiqin Chen1  Guanzhi Wang5  Xiusi Chen3 Wei Cheng4  Yisong Yue5  Ziniu Hu5 1

### v1

@@ 1 Introduction
Recent studies have demonstrated how Large Language Models (LLMs) can be utilized to learn skills for improved decision-making in interactive environments (Wang et al., 2023; 2022). However, learning skills in adversarial environments with multiple agents presents a significant challenge for LLMs, as it requires accounting for the responses of other players or environment to their actions. In these environments, it is hard to determine the optimal policy, since strategies that are effective against one opponent’s policy may not work for another, and the opponents can adapt to our current strategy. Moreover, the complexity of these environments means that we need to smartly search across a large policy space to find the optimal policy. Hence, existing LLM self-improvement methods that rely only on feedback from the LLM often have trouble distinguishing what improvements help and end up in local optima strategies that can be exploited.
[…]
Specifically, our approach combines LLM self-improvement with search, both on a high level strategy (or skill) space, and also on a low-level action space. On the high-level we do this by constructing a strategy tree in an evolutionary process, where we improve upon previously generated strategies. In our game-play settings, we specifically learn a value heuristic to evaluate game states, and a dialogue guide to guide the dialogue generation process. During the process, we also learn how to improve strategies, and what types of improvements tend to be more helpful generally. We search across the strategy tree using exploration and exploitation methods, eventually finding good strategies. No parameter updates or supervised fine-tuning are necessary during the search process.
The strategies are tested out through simulated self-play on the low level. By mastering a high-level strategy, the LLM not only learns a good policy for the player, but also gains insights into how opponents might react to that policy. With this information, our method can simulate the behaviors of both players and opponents by MCTS, gathering feedback through the simulated trajectories and the search process. In other words, we leverage the world-knowledge in the LLM to simulate behaviors of both players and opponents in order to get better feedback and plan better.
We apply our method to two games – Game of Pure Strategy (GOPS) and Resistance: Avalon – and show that our improvement process works for both games with dialogue and games without. We find that our improvement method is better than many other self-improvement methods, including both existing LLM-based self-improvement methods and traditional RL-based training approaches to AI in games, given the same computation and data budget. We also show the effectiveness of our method of collecting simulated self-play feedback, including using search to provide feedback.
To summarize, our main contributions are as follows:
  • We propose a general non-parametric LLM skill-learning framework, in which a low-level self-play tree-search provides value feedback, and a high-level skill learning module improves the skill in various forms, including functions and textual strategies.
  • To improve the sample efficiency of skill learning, we introduce a modular search method, which factorizes the process by an idea-implementation search with bandit exploration.
  • We apply this framework to GOPS and Avalon, and show that through learning a value heuristic and dialogue strategy guide we can learn a good policy for both action planning and dialogue generation. Some examples of good strategies learned through our method are shown in G.3 and H.3, which achieve higher win-rates against other agents.
2 Methodology
2.1 Strategy Learning in Decision Making Setting
[…]
The general goal in our decision-making setting is to learn a good policy function in a sequential decision-making setting (generally formulated as a partially observable Markov decision game (POMDG)), which can be done by improving strategies associated with the policy function. We describe in more detail what a strategy looks like, how we derive a policy function from a strategy, and how to acquire feedback for the strategy for both dialogue generation and action generation here.
Problem definition. Given state space $\mathcal{S}$ and action space $\mathcal{A}$, a policy function $\phi$ in policy space $\Phi$ is a mapping $\phi:\mathcal{S}\rightarrow\Delta\mathcal{A}$ where we allow $\phi$ to output a probability distribution over the actions ($\Delta\mathcal{A}$). An environment $\mathcal{E}=\langle\mathcal{S},\mathcal{A},\mathcal{N},T,R,A,\phi_{\epsilon}\rangle$ defines the state space $\mathcal{S}$, the action space $\mathcal{A}$, a set of actors $\mathcal{N}$, a transition function $T:\mathcal{S}\times\mathcal{A}\rightarrow\mathcal{S}$, a reward function $R:\mathcal{S}\times\mathcal{A}\rightarrow\mathbb{R}^{|\mathcal{N}|}$ that specifies intermediate rewards for each actor, an action function $A:\mathcal{S}\rightarrow\mathcal{N},\mathcal{P}(\mathcal{A})$ that specifies which actor may take what legal actions at some state where $\mathcal{P}$ is power set, and $\phi_{e}$, the policy function for the environment actor. Note that transitions are deterministic in our notation, and stochastic transitions are handled by the environment actor $e\in\mathcal{N}$ instead, so we cover stochastic, deterministic, multi-agent, and single agent cases. For a partial information setting, we also have a space of information sets $\mathcal{I}$ and a function $H:\mathcal{S}\times\mathcal{N}\rightarrow\mathcal{I}$ that maps from hidden states and actors to hidden information sets. Hence, $\phi:\mathcal{I}\rightarrow\Delta\mathcal{A}$ is a function from information sets to action distributions instead.
Let $f:\Sigma\rightarrow\Phi$ be the function that maps strategies to policies. A high-level strategy $\sigma\in\Sigma$ helps parameterize policies so that we can search over the lower dimension $\Sigma$ space instead of $\Phi$. Let $\Phi_{-i}$ denote the space of possible opponent policies. Then our goal is to find the optimal strategy $\sigma_{i}$ that approximates finding the optimal policy given the policies of the other agents $\phi_{-i}$, i.e.
where $\tau=(s_{0},a_{0},...)$ is the simulated trajectory according to the joint policy $(\phi_{i},\phi_{-i})$ and the transition function $T$, with $a_{t}\sim\phi(a_{t}|s_{t})$ and $s_{t+1}=T(s_{t},a_{t})$.
The state space, action space, and actor space are different depending on the setting. In non-stochastic, single agent settings such as question answering  (Yang et al., 2018), $\mathcal{N}=\{0\}$. In stochastic single agent settings such as WebShop  (Yao et al., 2022a), $\mathcal{N}=\{-1,0\}$ where we add an environment agent $-1$. We specifically focus on adversarial, multi-agent stochastic game ($\mathcal{N}=\{-1,0,1,...\}$) settings where the other agents are actively working against each other. In non-dialogue-based card games such as GOPS (see B for rules) for example, $\mathcal{S}$ consists of the cards played so far, $\mathcal{N}=\{-1,0,1\}$, and $\mathcal{A}$ consists of the cards you can play. In dialogue-based games such as Avalon (see A for rules), $\mathcal{A}$ consists of both the discrete actions (moves) such as voting, and also all possible natural language texts of a given length. Similarly, $\mathcal{S}$ consists of both the historical moves and historical dialogue record, and $|\mathcal{N}|\geq 6$.
Skills (strategies) here refer to any high-level strategy space $\Sigma$ where strategies $\sigma$ can be learned. In our Avalon game setting for example, we learn (1) a value heuristic $\sigma:=v:\mathcal{S}\rightarrow\mathbb{R}^{|\mathcal{N}|}$ (value function, see Appendix G.3 for example), a strategy to evaluate different states of the game (served as world model) and (2) a textual strategy guide on how to evaluate and analyze the current discussion and the steps for generating good dialogue (see Appendix H.3 for example).
We focus on learning high-level skills because this allows for higher level strategic planning and LLM reasoning, abstracting away low-level details that may be scenario specific and thus generalizing better to different cases. Given the size of the search space, it is also computationally expensive to search for the best action for each input state. While the skill coach helps us learn high-level strategies, the evaluator executes and evaluates the high-level strategies on the lower level. In our setting, this is usually done through game simulations and self-play, where the evaluator compares agents who use different strategies against each other.
@@ 2.1 Strategy Learning in Decision Making Setting
Figure 2: Overview of our skill coach improvement method. There are two improvement steps – the idea generation step and the strategy improvement step and we alternate between the two.
@@
During the idea generation step, a strategy $\sigma$ and its simulated trajectory feedback $\tau_{\sigma}$ is first selected from the strategy tree according to the adaptive selection policy. The feedback usually trajectories from previous self-play simulations, including the states visited, actions taken at those states, the estimated win-rate at that state, the final outcome of the trajectory, and any intermediate values used during simulations at those states. Since these trajectories are very long (around 30 steps each for Avalon), we select some key states in the trajectory to translate into natural language. We then prompt the LLM to reflect on this feedback and propose a few new ideas on how to improve the function. These new ideas are added to the idea queue, along with a score prior to how good the idea is.
@@
During the strategy improvement step, we first select a strategy and an idea from the libraries using the adaptive selection policy. We then prompt the LLM to implement the idea of the strategy, generating a new improved strategy. We then evaluate the strategy using the evaluator, which conducts self-play simulations with the strategy, and records the simulated trajectory data. During simulations, players conduct an MCTS tree search to estimate the expected win-rate at different states, which provides additional feedback. We add the improved strategy (and its performance) to the strategy tree, and update the improvement score for the idea that was used.
@@
A key challenge during the improvement process is that it is often difficult to evaluate whether an improved strategy is better than the old one and why. Simply querying the LLM to improve the strategy based on the feedback without guidance from an improvement idea often results in the LLM making changes to all parts of the strategy. Thus it is hard to determine what specific changes affected the strategy after we evaluate it. Hence, we utilize an additional idea queue to modularize the search process. Instead of improving everything all at once, we incrementally improve the strategy, ceteris paribus, creating a more scientific testing and improvement process that avoids confounding factors.
@@
Moreover, the idea queue helps us keep track of what improvements have been successful in the past, which are also likely to help in the future. Since the strategies are similar and used in the same setting, improvements from one can be transferred to another. These improvements, often involving additive adjustments like penalties or extra questions, are generalizable across strategies. We also specifically prompt for high-level, generalizable ideas. For example, in the case of improving a dialogue guide, an improvement idea might be to “Add a question asking the player to assess their level of suspicion towards each player and craft a statement addressing how they will interact with those players during the discussion phase.” In the case of value heuristics generation, an example might be “Revise the dynamic penalty adjustment to more accurately reflect the impact of high-value score cards left in the deck.” Both of these ideas are general enough to be applicable to different strategies.
@@
To make the improvement process more manageable, we optimize the policy for dialogue generation and moves separately in Avalon. We then integrate the two parts (refer to Appendix D), which together produces an agent fully capable of playing games with discussion. This type of integration of separate components has seen success in other settings such as Diplomacy  († et al.(2022)(FAIR)†, Bakhtin, Brown, Dinan, Farina, Flaherty, Fried, Goff, Gray, Hu, et al., FAIR).
2.3 Self-improvement Feedback for Non-dialogue Actions
For non-dialogue actions, while the action spaces $\mathcal{A}$ and state spaces $\mathcal{S}$ themselves are usually discrete and finite, the number of possible functions $\Phi$ from state space to action space is very large. Most LLM-agents query the LLM directly with the state information for next actions in decision-making environments  (Yao et al., 2023; Shinn et al., 2024; Zhao et al., 2024). However, we found this method to be costly since the LLM needs be queried for every move, and a game of Avalon usually has at least 20 moves per player. This becomes even costlier when we use look-ahead search to improve the planning capabilities of the agent and does also needs to query the LLM for every look-ahead future actions and states. Traditionally reinforcement learning tackles the problem of having a large policy space by parameterizing the policy and optimizing the parameters instead, starting from a random initialization of the parameters, thus reducing the search space. Building on this approach, instead of querying the LLM every time for an action, we instead parameterize the policy as a python function, and we have the LLM search over the space of such functions instead.
With its inherent world model, the LLM can help us search and optimize over the policy space more effectively. Given the rules of the game in natural language form that a human can understand, the LLM can quickly generate policies that seem reasonable. In natural language form, it is often easier to describe the value of a state versus describing the optimal action to take in a given state. Humans often mentally go through the process of comparing different states (outcomes) when making decisions. Thus, we opt to learn a value function, i.e. $\sigma:=v:\mathcal{S}\rightarrow\mathbb{R}^{|\mathcal{N}|}$, an approach that is often used in reinforcement learning to help stabilize the training process. It is easy to convert from a value function to a policy function since we can simple take the action that leads to the best state, i.e. $\phi_{i}(s)=\mathop{argmax}_{a\in\mathcal{A}}Q(s,a)=\mathop{argmax}_{a\in\mathcal{A}}R_{i}(s,a)+v_{i}(s^{\prime})|s^{\prime}=T(s,a)$.
We call it a value heuristics function because, given the little data that we improve on, it is an inaccurate estimate of the true value function. Hence, to resolve these inaccuracies, we additionally enhance the policy function with MCTS, used commonly in other AI agents such as Alpha-go and Muzero  (Silver et al., 2017; Schrittwieser et al., 2020). The MCTS process is depicted in Figure 10. We provide more details on our MCTS implementation in the Appendix E.
The search process allows us to get better value estimates than the initial ones given by the value heuristic when making decisions. More importantly, this search process also allows us to generate more feedback than normally possible, since we can compare the updated value estimate computed through MCTS with the initial estimate from the value heuristic. The estimated win-rate from search also provides a shaped process reward, which is better than the outcome reward of either winning or losing as a reward signal. Specifically, we select key states that have the maximum discrepancy between the MCTS estimate and the outcome of the episode to provide as feedback for self-improvement.
2.4 Self-improvement Feedback for Dialogue Generation
Dialogue generation presents another distinct challenge for self-improvement. In dialogue generation, both the action spac $\mathcal{A}$, i.e. the number of possible sequences of words one could generate for any discussion round, the state space $\mathcal{S}$, i.e. the number of possible input dialogue from previous rounds, are huge. This means that the number of possible dialogue generation policies $\Phi$ is massive, and traditional parameter training approaches have great difficulty optimizing across this space. We solve this problem by learning a high-level strategy guide that the agent uses when speaking. Specifically, the strategy guide formalizes a process to think about how to generate dialogue, given the situation. This is implemented as a question and answer worksheet. The agent is first instructed to answer all the questions in the strategy guide before using it as a prompt to generate dialogue.
Example LLM Generated Value Heuristic Function Moreover, it is difficult to provide good reward signals during the dialogue training process. The first problem is the lack of available training data. Existing dialogue generation methods for games often rely on imitation learning on existing human generated text data through parameter training of the underlying network  († et al.(2022)(FAIR)†, Bakhtin, Brown, Dinan, Farina, Flaherty, Fried, Goff, Gray, Hu, et al., FAIR). However, in many settings, such large quantities of human generated data might not be readily available for the tuning of language models. The second problem lies in how to accurately evaluate how the generated dialogue performs. Players need to optimize and balance multiple objectives when discussing. For example, in Avalon, the Merlin player needs to both hide themselves from Evil and convince the Good players that they are Good through discussion. Example Dialogue Strategy Guide Q1: Who are the players you suspect are on the Evil side and why? A1: I suspect Player 2 and Player 4 are on the Evil side. Player 2’s suggestion for the team consisting of themselves and Player 1 seems too convenient, as it would allow Evil to easily fail the quest. Player 4’s agreement with Player 1 and Player 2’s proposed team without offering any alternative suggestions raises suspicion. Q2: Draft a subtle statement that casts doubt on these players without directly accusing them. A2: "I can’t help but wonder if some players are pushing for certain team compositions to benefit themselves rather than considering what is truly best for the success of our Quest." Generated dialogue: “I have observed patterns in the voting that suggest certain players may be working together to push for specific team compositions. This raises suspicion and calls into question their true intentions. We must remain vigilant and not allow ourselves to be swayed by potentially deceitful tactics.” We solve these problems by first simulating a game using some initial policy $\phi_{0}$ and storing the simulation in a scenario database. Then during evaluation, we sample scenarios from the database to evaluate with. A scenario is a state in the game when a player had to take a dialogue action. We give the player a history of the previous discussions and actions up to the decision point as if it played the game up to that point, and ask it to generate new dialogue using the new strategy guide. Then we simulate the thought process of the opponents as if they also played the game up to that point, and we get their belief updates. Using the belief updates we then score the performance of the player that uses the improved strategy guide, and provide the opponent thought process as feedback. We provide more details on our implementation in Appendix F.

### v2

@@ 1 Introduction
Recent studies have shown the potential of Large Language Models (LLMs) for learning skills and improving decision-making in interactive environments Wang et al. (2022; 2023a; 2024); Xi et al. (2023); Zhao et al. (2024); Liu et al. (2023). However, adversarial, multi-agent environments present a significant challenge. LLMs must reason about opponent actions, plan ahead strategically, and navigate a vast policy space Kambhampati et al. (2024); Valmeekam et al. (2023); Kambhampati (2024). This complexity hinders the LLM’s ability to identify, understand, and translate optimal policies into effective actions. Simply generating decision rules by querying the LLM proves inefficient and impractical in such settings. This raises two critical questions: (1) How can LLMs effectively learn complex policies in adversarial multi-agent environments? (2) What is the best approach to improve these policies?
@@
To summarize, our main contributions are as follows:
[…]
  • We introduce a modular method of improving high level strategies using population based self-play, without any training data. We demonstrate through experiments that our improvement method outperforms other LLM improvement and skill learning methods.
  • We apply this framework to GOPS and Avalon, complex competitive multi-turn games, and show how we can learn good policies for both dialogue and non-dialogue actions, outperforming existing agents. Some examples of good strategies learned through our method are shown in H.3 and I.3, which achieve higher win-rates against other agents.
2 Methodology
We cover the problem formulation, how we create high level strategy abstractions, and how we improve strategies using LLMs and self-play in the following sections.
2.1 Decision making setting and games
The general goal in our decision-making setting is to learn a good policy function in a sequential decision-making setting (generally formulated as a partially observable Markov decision game (POMDG)). Our strategy-learning framework is generalizable to any decision making setting.
Problem definition. Given state space $\mathcal{S}$ and action space $\mathcal{A}$, a policy function $\phi$ in policy space $\Phi$ is a mapping $\phi:\mathcal{S}\rightarrow\Delta\mathcal{A}$ where we allow $\phi$ to output a probability distribution over the actions ($\Delta\mathcal{A}$). An environment $\mathcal{E}=\langle\mathcal{S},\mathcal{A},\mathcal{N},T,R,A,\phi_{\epsilon}\rangle$ defines the state space $\mathcal{S}$, the action space $\mathcal{A}$, a set of actors $\mathcal{N}$, a transition function $T:\mathcal{S}\times\mathcal{A}\rightarrow\mathcal{S}$, a reward function $R:\mathcal{S}\times\mathcal{A}\rightarrow\mathbb{R}^{|\mathcal{N}|}$ that specifies intermediate rewards for each actor, an action function $A:\mathcal{S}\rightarrow\mathcal{N},\mathcal{P}(\mathcal{A})$ that specifies which actor may take what legal actions at some state where $\mathcal{P}$ is power set, and $\phi_{\epsilon}$, the policy function for the environment actor. Note that transitions are deterministic in our notation, and stochastic transitions are handled by the environment actor $\epsilon\in\mathcal{N}$ instead, so we cover stochastic, deterministic, multi-agent, and single agent cases. For a partial information setting, we also have a space of information sets $\mathcal{I}$ and a function $H:\mathcal{S}\times\mathcal{N}\rightarrow\mathcal{I}$ that maps from hidden states and actors to hidden information sets. Hence, $\phi:\mathcal{I}\rightarrow\Delta\mathcal{A}$ is a function from information sets to action distributions instead. A strategic profile $\boldsymbol{\phi}=[\phi_{i}]_{|\mathcal{N}|}$ describes the policies for every player $i\in\mathcal{N}$.
Let $f:\Sigma\rightarrow\Phi$ be the function that maps strategies to policies. A high-level strategy $\sigma\in\Sigma$ helps parameterize policies so that we can search over the lower dimension $\Sigma$ space instead of $\Phi$. Let $\Phi_{-i}$ denote the space of possible opponent policies, where $-i$ are the indices of players other than i. Then our goal is to find the optimal strategy $\sigma_{i}$ that approximates finding the optimal policy given the policies of the other agents $\phi_{-i}$, i.e.
where $\tau=(s_{0},a_{0},...)$ is the simulated trajectory according to the strategic profile $(\phi_{i},\phi_{-i})$ and the transition function $T$, with $a_{t}\sim\phi(a_{t}|s_{t})$ and $s_{t+1}=T(s_{t},a_{t})$. Thus, a key part of these settings involves learning the probable strategic profile of other players, usually by assuming that they play optimally like in a Nash equilibrium  (Perolat et al., 2022).
The state space, action space, and actor space are different depending on the setting. In non-stochastic, single agent settings such as question answering  (Yang et al., 2018; Wang et al., 2022; Shridhar et al., 2020; Thorne et al., 2018), $\mathcal{N}=\{0\}$. In stochastic single agent settings such as web browsing  (Yao et al., 2022a; Deng et al., 2024; Zhou et al., 2023), $\mathcal{N}=\{\epsilon,0\}$ where we add an environment agent $\epsilon$. We specifically focus on adversarial, multi-agent stochastic game ($\mathcal{N}=\{\epsilon,0,1,...\}$) settings where the other agents are actively working against each other. In non-dialogue-based card games such as GOPS (see B for rules) for example, $\mathcal{S}$ consists of the cards played so far, $\mathcal{N}=\{\epsilon,0,1\}$ for players 1 and 2 respectively, and $\mathcal{A}$ consists of the cards you can play. In dialogue-based games such as Avalon (see A for rules), $\mathcal{A}$ consists of both the discrete actions (moves) such as voting, and also all possible natural language texts of a given length. Similarly, $\mathcal{S}$ consists of both the historical moves and historical dialogue record, and $\mathcal{N}=\{\epsilon,0,1,...\}$ depending on the number of players involved, with $|\mathcal{N}|\geq 6$.
2.2 Abstracting high level strategies as processes
A key part of our framework abstracts key features of the policy $\phi$, representing them as a high-level strategy $\sigma$, which is more suitable for LLM-driven improvement. $\sigma$ is then executed and refined by a low-level executor, often during inference, resulting in the execution policy $\phi_{\sigma}$.
[…]
Abstraction offers two key benefits: (1) High-level strategies provide a more abstract, compressed, and intuitive representation of the problem, which the LLM better understands, enabling its generalization and reasoning capabilities. (2) Searching over high-level strategies is more efficient than low-level policy exploration, as it simplifies the search space by focusing on core principles rather than fine details. This approach guides the search toward promising areas more quickly, leveraging domain knowledge, patterns, or heuristics that generalize across scenarios.
For non-dialogue actions, while the action spaces $\mathcal{A}$ and state spaces $\mathcal{S}$ are typically discrete and finite, the number of possible functions $\Phi$ mapping state to action is vast. Most LLM-agents query the LLM directly with state information to decide the next action in decision-making environments  (Yao et al., 2023; Shinn et al., 2024; Zhao et al., 2024). However, this is costly because the LLM must be queried for every move, and in Avalon, each player makes at least 20 moves. This becomes even more expensive when incorporating look-ahead search, which requires querying the LLM for future actions and states. Traditionally, policy-gradient reinforcement learning addresses the large policy space by parameterizing the policy and optimizing these parameters, reducing the search space.
We abstract and “parameterize” the policy $\phi$ as a value heuristic (VH) $\sigma:=v:\mathcal{S}\rightarrow\mathbb{R}^{|\mathcal{N}|}$ that estimates the expected cumulative returns for each player at a given state, typically the expected probability of winning for each player. Using a value heuristic simplifies reasoning, as it’s easier to describe how good a state is than to specify the optimal action. This makes it intuitive for the LLM to reason about winning probabilities. Additionally, we represent the value heuristic as Python code, allowing the LLM to easily process and modify the logic for computing the value for each state.
Example LLM Generated Value Heuristic Function At the low-level, we execute the strategy (value function) by selecting the action that leads to the best state: $\phi_{i}(s)=\mathop{argmax}_{a\in\mathcal{A}}Q(s,a)=\mathop{argmax}_{a\in\mathcal{A}}R_{i}(s,a)+v_{i}(s^{\prime})|s^{\prime}=T(s,a)$. Since we learn the value $v_{i}$ for each actor, we can compute the best action for each actor $i$, refining the value function into a strategic profile $\boldsymbol{\phi}$. Since value heuristics can be inaccurate, we refine the policy using MCTS to look ahead through multiple action-state sequences, resulting in a better policy  (Grill et al., 2020; Silver et al., 2017; Schrittwieser et al., 2020). MCTS also generates additional feedback by comparing the updated value estimate from MCTS with the initial value heuristic.The estimated win-rate from search provides a shaped reward signal, which is more informative than the simple win/lose outcome reward. We focus on key states with the largest discrepancy between the MCTS estimate and the episode outcome for self-improvement feedback. More details on our MCTS implementation are in Appendix F. Dialogue generation poses a unique challenge for self-improvement. Both the action space $\mathcal{A}$, representing possible word sequences for each discussion round, and the state space $\mathcal{S}$, representing prior input dialogues, are enormous. As a result, the number of possible dialogue generation policies $\Phi$ is vast, making traditional parameter training approaches difficult to optimize. To address this, we develop a high-level strategy guide that the agent processes before speaking. This guide formalizes a chain-of-thought process to analyze the situation before generating dialogue  (Wei et al., 2022b). It is implemented as a question-and-answer worksheet, where the agent answers all questions in the guide before using it as a prompt for dialogue generation. Example (fully generated) Chain of Thought Dialogue Strategy Guide Q1: Who are the players you suspect are on the Evil side and why? A1: I suspect Player 2 and Player 4 are on the Evil side. Player 2’s suggestion for the team consisting of themselves and Player 1 seems too convenient, as it would allow Evil to easily fail the quest. Player 4’s agreement with Player 1 and Player 2’s proposed team without offering any alternative suggestions raises suspicion. Q2: Draft a subtle statement that casts doubt on these players without directly accusing them. A2: "I can’t help but wonder if some players are pushing for certain team compositions to benefit themselves rather than considering what is truly best for the success of our Quest." Generated dialogue: “I have observed patterns in the voting that suggest certain players may be working together to push for specific team compositions. This raises suspicion and calls into question their true intentions. We must remain vigilant and not allow ourselves to be swayed by potentially deceitful tactics.”
2.3 Modular search for high-Level strategy improvement
The framework for our high-level strategy improvement process is shown in Figure 2 with corresponding pseudo-code in 1 and more detailed diagram in Figure 9. Each improvement cycle has two steps: (1) reflection and idea generation, and (2) strategy improvement. In the idea generation step, we prompt the LLM to reflect on simulated self-play feedback from previous strategies and generate improvement ideas, which are added to the idea queue. In the strategy improvement step, we select a strategy from the strategy tree and an idea from the queue, prompting the LLM to refine the strategy. The improved strategy is then evaluated through self-play simulations, with feedback and reward signals guiding future improvements.
Figure 2: Overview of our high-level strategy improvement method: The process alternates between two steps—idea generation and strategy improvement. For more details see Figure 9.
The improvement process maintains a strategy library, starting with seed strategies, along with each strategy’s performance score and raw feedback. It also keeps a queue of improvement ideas, tracking how much each idea improves strategies on average (score). Strategies are connected in a tree structure, showing how they evolved and which improvement ideas were used. The process alternates between two loops: the reflection and idea generation step and the strategy improvement step, running for a fixed number of cycles. An adaptive selection policy, such as UCB or BFS, is used to choose strategies and ideas for improvement.
2.3.1 Idea Generation
@@
A sampled strategy $\sigma$ and its simulated feedback trajectory $\tau_{\sigma}$ are selected from the strategy tree using the adaptive selection policy. The feedback consists of trajectories from previous self-play simulations, including visited states, actions taken, estimated win-rates, final outcomes, and intermediate values. Since these trajectories can be lengthy, we select key states to translate into natural language. The LLM is then prompted to reflect on this feedback and propose new improvement ideas, which are added to the idea queue with a prior score estimating their potential effectiveness.
[…]
2.3.2 Strategy Improvement
@@
We first sample a strategy and an idea from the libraries using a method that balances exploration and exploitation such as UCB. We then prompt the LLM to implement the idea of the strategy, generating a new improved strategy. We then evaluate the strategy using the evaluator, which conducts self-play simulations with the strategy, and records the simulated trajectory data. During simulations, players conduct an MCTS tree search to estimate the expected win-rate at different states, which provides additional feedback. We add the improved strategy (and its performance) to the strategy tree, and update the improvement score for the idea that was used.
[…]
2.3.3 Search Modularization
@@
A key challenge in the improvement process is evaluating whether a new strategy is better than the previous one and understanding why. Simply prompting the LLM to improve based on feedback without a clear idea often leads to changes across all parts of the strategy, making it difficult to pinpoint which changes had an impact. To address this, we use an idea queue to modularize the search process. Instead of improving everything at once, we incrementally refine the strategy, ceteris paribus, enabling a more scientific approach that avoids confounding factors.
@@
The idea queue also helps track successful improvements, which are likely to be useful in the future. Since strategies are similar and used in the same context, improvements from one strategy can be transferred to another. These improvements, often involving additive adjustments like penalties or extra questions, are generalizable across strategies. We also prompt for high-level, generalizable ideas. For example, in a dialogue guide, an idea might be to add a certain question. For value heuristics generation, an idea might be to add a dynamic penalty term based on high-value cards. These general ideas apply to various strategies. Empirically, we found that transferable and additive improvements boost performance, as shown in Table 1.
@@
To make the improvement process more manageable, we optimize the policy for dialogue generation and moves separately in Avalon. We then integrate the two parts (refer to Appendix E), which together produces an agent fully capable of playing games with discussion. This type of integration of separate components has seen success in other settings such as Diplomacy † et al.(2022)(FAIR)†, Bakhtin, Brown, Dinan, Farina, Flaherty, Fried, Goff, Gray, Hu, et al. (FAIR).
2.4 Population based self-play simulation
Recent works focus on using either another LLM for critique-based feedback  (Madaan et al., 2024; Shinn et al., 2024), real environment interactions  (Nottingham et al., 2024), or a combination of both  (Wang et al., 2023a) during the LLM-improvement process. Our method improves on this by simulating opponents using the learned high-level strategy, enabling higher-quality feedback. Since our high-level strategies can be refined into a policy for any player, we can easily pit different strategies against each other by refining them into specific policies for different players. Specifically, we use population based self-play, where we conduct round-robin games among the top ten strategies generated and use the average performance as the strategy score for each strategy. During round-robin games, strategies are given the same level of low-level refinement to ensure fairness. This evolutionary approach makes the final improved strategy more robust to different opponent policies, since it will have survived multiple round-robin tournaments. We demonstrate the effectiveness of this method in section 3.4.